OpenStudy (anonymous):

Let f(x) = x2 + x + 2 and g(x) = 2x2 + 5. Find f(g(x)). Show each step of your work. I put them together for (2x2+5)2+( 2x2+5)+2. Now I'm stuck.

OpenStudy (firekat97):

I think you just simplify now, remember for \[(a + b)^2 = a^2 + 2ab + b^2\]

OpenStudy (firekat97):

so try and simplify and post up what you get

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

I'm still confused. the exponents are what I don't understand. Do I add the terms together or...

OpenStudy (firekat97):

yup you want to add like terms together. have you tried expanding \[(2x^2 + 5)^2\]?

OpenStudy (anonymous):

Expanding?

OpenStudy (firekat97):

because it best if you have a go and I can then check and direct you

OpenStudy (firekat97):

Yup you can expand/open them using the rule I'd shown you above

OpenStudy (firekat97):

then we can collect like terms to simplify

OpenStudy (anonymous):

Oh! Ok. Got it. Hold on a moment.

OpenStudy (anonymous):

I got \[2x ^{4}+5^{2}+2x ^{2}+7\] Is that all I can do? Or can it be simplified more?

OpenStudy (firekat97):

hm your answer doesn't quite match mine, what did you get when you expanded \[(2x^2 + 5)^2\]?

OpenStudy (anonymous):

\[2x ^{4}+5^{2}\]

OpenStudy (anonymous):

OH!

OpenStudy (anonymous):

5squared should be 25.

OpenStudy (firekat97):

not quite, so we have \[(2x^2 + 5)^2\] and using \[(a + b)^2 = a^2 + 2bc + b^2\] we should get \[(2x^2)^2 + 2(2x^2)(5) + 5^2\] which simplifies down to \[4x^4 + 20x^2 + 25\] when we expand

OpenStudy (anonymous):

Is That all we can do?

OpenStudy (firekat97):

no, we'e simplified the \[(2x^2 + 5)^2\] part, but don't forget, our task is to simplify \[(2x^2 + 5) + (2x^2 + 5) + 2\] so now we have \[4x^4 + 20x^2 + 25 + (2x^2 + 5) + 2\] so now you gotta collect like terms, try this and show me what you get @Whitelynx

OpenStudy (firekat97):

sorry correction from above**** "Our task is to simplify \[(2x^2 + 5)^2 + (2x + 5) + 2\]"****

OpenStudy (anonymous):

\[4x ^{4}+22x ^{2}+32\] Is this right?

OpenStudy (firekat97):

yupp thats correct @whitelynx! :D